OpenStudy (anonymous):

(x+5)2 =-17 =5x*2=10x =10/17

OpenStudy (anonymous):

not sure of your question

Directrix (directrix):

Would you check to see if you wrote the problem correctly? Is is a system of 2 equations?

OpenStudy (mimi_x3):

Is it: \[(x+5)^2 = -17x\]? or \[2(x+5)=-17\]

OpenStudy (mimi_x3):

Woops, typo \[(x+5)^2=-17\]